need help asap!!!A 23 kg body is moving through space in the positive direction of an x axis with a speed of 340 m/s when, due to an internal
Answer #1

Solatn) 2 U 340Ms n+ x-asis U= 340 () mls M si mls in 1 dielion, U+1) 00 msin kdivectin, 2-oo -) 23-(4.9+54) 12.k / M3=(m) -(